Seven years ago, my father's age was five times mine. Right now, his age is the sum of our two ages 7 years ago. How old am I? Thanks for your help.
Answer by TimothyLamb(4379)

x = father age now
y = son age now
---
x - 7 = 5(y - 7)
x = (x - 7) + (y - 7)
---
put the system of linear equations into standard form
---
x - 7 = 5y - 35
x = x - 7 + y - 7
---
x - 5y = - 28
y = 14
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = father age now = 42
y = son age now = 14
---
now by substitution:
---
x - 5y = - 28
y = 14
---
x - 5*14 = - 28
x = 5*14 - 28
x = 42
---
solution:
x = father age now = 42
y = son age now = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php